OpenStudy (anonymous):

please show steps: Solve: log base 5 3-1/2 log base 5 x=log base 5 1/2

OpenStudy (anonymous):

0.0 ! how ? did u get the sqare root of 3

hero (hero):

assuming that the negative one-half is an exponent. Is it?

OpenStudy (anonymous):

i don't know... this si how it is written \[\log_{5} 3-1/2\log_{5} x =\log_{5} 1/2\]

hero (hero):

I see.

OpenStudy (anonymous):

log base 5 3-1/2 log base 5 x=log base 5 1/2 rearrange to have X as subject. 1/2log base 5 of x =log base 5 of 3 - log base 5 of 1/2 => log base 5 of x^1/2=log base 5 of 3/[1/2] => log base 5 of x^1/2=log base 5 of 6 take anti logs on both sides ( since they are all logs of the same base) => x^1/2 = 6 => x= 6^2 = 36

OpenStudy (anonymous):

FINALLY thank you

OpenStudy (anonymous):

welcome The trick perhaps is logx -logy = log(x/y) as long as they are all of the same base

hero (hero):

izzy, had you posted it clearly the first time, you would have gotten a solution much quicker. The whole time, people were trying to decipher what you had written initially.

OpenStudy (anonymous):

i posted as written, thank you very much